IMPROVEMENT OF SPECTROSCOPIC CONSTANTS FOR THE A^{3}Π_{1u} ← X^{1}Σ_{g}^{+} SYSTEM OF Br_{2}

\underline{NOBUO NISHIMIYA}, TOKIO YUKIYA, and MASAO SUZUKI, Department of Electronics and Information Technology, Tokyo Polytechnic University, Iiyama 1583, Atsugi-City, 243-0297 Kanagawa, Japan; ROBERT J. LE ROY, Department of Chemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

High-resolution spectroscopy in the near infrared region is very important for frequency standards. We have measured the Doppler limited vibrational-rotational spectrum of ICl,^{a} IBr,^{b} I_{2}^{c} and Br_{2}^{d} by using a titanium sapphire laser, and have reported their spectroscopic constants. The present work presents revised spectroscopic constants for the X-state of Br_{2} which are obtained by performing a combined-isotopologue least-squares fit to all available data for the three isotopic species. We are also attempting to observe the lowest vibrational levels of the A-state in order to allow us to determine the potential energy function of this state. Measurements were made at room temperature in the region 13000 − 13700 cm^{−1} and at 150-250^{∘}C in the region 11600 − 13000 cm^{−1}. Details of these results will be presented.

a N. Nishimiya, T. Yukiya, and M. Suzuki, J. Mol. Spectrosc. 163, 43 (2009).

bT. Yukiya, N. Nishimiya, and M. Suzuki, J. Mol. Spectrosc. 214, 132 (2002).

cN. Nishimiya, T. Yukiya, and M. Suzuki, J. Mol. Spectrosc. 182, 271 (1997).

dT. Yukiya, N. Nishimiya, and M. Suzuki, Columbus Meeting, paper WH02 (2005).